OpenStudy (anonymous):

find the geometric mean of 4 and 25

OpenStudy (istim):

There a difference between regular mean and geometric mean?

OpenStudy (anonymous):

Yes, geometric mean of 2 numbers is the square root of the product of the numbers

OpenStudy (anonymous):

so 10 in this case

OpenStudy (mertsj):



OpenStudy (anonymous):

\[\sqrt{25*4}=\sqrt{100}=10\]

OpenStudy (mertsj):

x is the geometric mean. Solve for it and you will see that rick is right.